The electric motor of a model train accelerates the train from rest to 0.620 m/s in 21.0cs. The total mass of the train is 875 g. Find the average power delivered to the train during the acceleration


1
Expert's answer
2021-11-07T19:28:05-0500

The average power is speed times force:


"P=Fv_\text{av}=mav_\text{av}=m\frac{v_\text{final}}{t}·\frac{v_\text{final}}{2}=\frac{mv_\text{final}^2}{2t},\\\space\\
P=8\text{ mW}."